OpenStudy (anonymous):

Given the linear equation ,y=-2/7x+3 find the y-coordinates of the points (-7, ), (0, ), and (7, ). Please show all your work. Plot those points and graph the linear .

OpenStudy (anonymous):

you given x=-7, and x=7 and x=7 put the x value in your equation & find y :)

OpenStudy (anonymous):

omg this is so hard!!!

OpenStudy (anonymous):

its not... try it! y=-2/7 *(-7) +3=2+3=...?

OpenStudy (anonymous):

sorry inik go ahead :)

OpenStudy (anonymous):

want them to do it... at least to try!

OpenStudy (anonymous):

ima try it and post my answer!

OpenStudy (anonymous):

sorry... I don't do answers only. it has to be some effort on your side....

OpenStudy (anonymous):

i want to be able to learn how to do it myself!

OpenStudy (anonymous):

and I want to help!

OpenStudy (anonymous):

-4.28

OpenStudy (anonymous):

The easiest way to do this is to plug in the x coordinates given into the equation. Then, you can get the y coordinate.

OpenStudy (anonymous):

let's try x=0: y= -2/7 *0 +3=...? it will give you point (0,...)

OpenStudy (anonymous):

the first one was: y=-2/7 *(-7) +3=2+3=...? =5 , right? so, the first coordinate: (-7,5) what is the second?

OpenStudy (anonymous):

(0,5)

OpenStudy (anonymous):

no... if x=0: y= -2/7 *0 +3=3 so your coordinate: (0,3) see it?

OpenStudy (anonymous):

let's try the last one: x=7, so y= -2/7 *7 +3=-2+3=... what is your last coordinate?

OpenStudy (anonymous):

yeah i think!

OpenStudy (anonymous):

(5,-7)